# FlowDreamer: Exploring High Fidelity Text-to-3D Generation Via Rectified Flow

### v2

@@ Abstract
Recent advances in text-to-3D generation have made significant progress. In particular, with the pretrained diffusion models, existing methods predominantly use Score Distillation Sampling (SDS) to train 3D models such as Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3D GS). However, a hurdle is that they often encounter difficulties with over-smoothing textures and over-saturating colors. The rectified flow model – which utilizes a simple ordinary differential equation (ODE) to represent a linear trajectory – shows promise as an alternative prior to text-to-3D generation. It learns a time-independent vector field, thereby reducing the ambiguity in 3D model update gradients that are calculated using time-dependent scores in the SDS framework. In light of this, we first develop a mathematical analysis to seamlessly integrate SDS with rectified flow model, paving the way for our initial framework known as Vector Field Distillation Sampling (VFDS). However, empirical findings indicate that VFDS still results in over-smoothing outcomes. Therefore, we analyze the grounding reasons for such a failure from the perspective of ODE trajectories. On top, we propose a novel framework, named FlowDreamer, which yields high-fidelity results with richer textual details and faster convergence.The key insight is to leverage the coupling and reversible properties of the rectified flow model to search for the corresponding noise, rather than using randomly sampled noise as in VFDS. Accordingly, we introduce a novel Unique Couple Matching (UCM) loss, which guides the 3D model to optimize along the same trajectory. Our FlowDreamer is superior in its flexibility to be applied to both NeRF and 3D GS. Extensive experiments demonstrate the high-fidelity outcomes and accelerated convergence of FlowDreamer. Moreover, we highlight the intriguing open questions, such as initialization challenges in NeRF and sampling techniques, to benefit the research community. For more details, please visit our project page at https://vlislab22.github.io/FlowDreamer.
@@ 1 Introduction
Recently, with the advances of text-to-2D image synthesis techniques based on the diffusion models, text-to-3D generation also undergoes a surge of research interest. A seminal work, DreamFusion (Poole et al., 2022) sets a cornerstone by proposing Score Distillation Sampling (SDS) to address this issue by leveraging pretrained text-to-image diffusion model, to train Neural Radiance Fields (NeRF)Mildenhall et al. (2021). It has been rapidly evolved to 3D Gaussian Splatting (3D GS) (Kerbl et al., 2023; Tang et al., 2023; Yi et al., 2023) for faster training and rendering.
Despite the success, some works, e.g., Lin et al. (2023); Poole et al. (2022); Zhu et al. (2023) unveil that SDS suffers from issues, such as over-smoothing textures and over-saturating colors. For this reason, some attempts, e.g.,Wang et al. (2024a); Liang et al. (2023); Wu et al. (2024) improve SDS from different perspectives. For example, Prolificdreamer (Wang et al., 2024a) introduces variational score distillation (VSD), which models 3D parameters as random variables to distill 3D assets, however, it requires significantly more time to optimize. Consistent3D (Wu et al., 2024) designs a consistency distillation sampling method to train the 3D model. Nevertheless, the quality improvements are still limited. Luciddreamer (Liang et al., 2023) proposes interval score matching (ISM) loss in the diffusing trajectory, but the loss drops many terms with the same scale.
Recently, flow matching approaches (Liu et al., 2022; Lipman et al., 2022) pave new ways for fast and high-quality generation. Among them, rectified flow model (Liu et al., 2022; Esser et al., 2024; Liu et al., 2023) uses a simple ordinary differential equation (ODE) to represent a straight trajectory. It trains a time-independent vector field, indicating that the direction of the vector field remains constant. Whereas diffusion trajectories result in curved trajectories(Lipman et al., 2022), score(Song et al., 2020b) is time-dependent, meaning that SDS optimizing over uniformly sampled values of $t$ can produce different gradient directions. Owning to its merits, we interestingly find that it could serve as an alternative prior for text-to-3D generation.
@@
In light of this, we first develop a mathematical analysis to seamlessly integrate SDS with rectified flow model. This enables us to build up an initial framework, named as Vector Field Distillation Sampling (VFDS). However, empirical results demonstrate that VFDS still leads to over-smoothing textures, see Figure 2. To this end, we further analyze the grounding reasons for such a failure from the perspective of ODE trajectories. This way, we find that, as VFDS randomly samples noise, it leads to multiple ODE trajectories in nearly the same images, i.e., from camera poses with mild differences, see Figure 4. This causes inconsistent update directions, leading to over-smoothing textural issues.
@@ 3.1 VFDS: SDS in the Lens of Rectified flow
Adapting SDS to the rectified flow framework. We first briefly introduce the rectified flow (Liu et al., 2022). Let $\pi_{1}$ and $\pi_{0}$ denote Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ and data distribution, respectively. $\epsilon$ and $x_{0}$ are respectively sampled from $\pi_{1}$ and $\pi_{0}$. Rectified flow defines the forward process as:
@@
Figure 3: (a): An illustration of the reversible and coupling properties of the rectified flow model. The coupling property indicates that $\epsilon$ is connected to $x$ by a single trajectory and does not cross other trajectories. The rectified flow trajectory is bidirectional, indicating that it is reversible. (b) An illustration of the UCM loss. Given a rendered image, the push-backward process is used to search for #$\#_{\phi}[x]$, followed by interpolation and then gradient calculation.
@@
where $w(t)$ is a time-dependent weighting function, $U[0,1]$ denotes the uniform distribution within $[0,1]$. Because the rectified flow model is an ODE model, it has reversible and coupling properties. The $\epsilon$ from the Gaussian noise distribution is uniquely coupled with the $x$ from the data distribution. Moreover, the rectified flow trajectory is bidirectional, indicating that it is reversible, as shown in Figure 3(a). Specifically,
@@
where $\theta$ is the 3D model parameters, $x=g(\mathbf{\theta,c})$ denotes a rendered image from a camera pose $\mathbf{c}$, $\epsilon$ denotes randomly sampled Gaussian noise. Following the convention of the SDS, we omit the transformer Jacobian term for effective training. Therefore, $\left(\frac{\partial v_{\phi}(x_{t},t)}{\partial x_{t}}\frac{\partial x_{t}}{\partial x}+1\right)$ becomes a constant and can be absorbed by $w(t)$, so we have
@@
Based on Equation 5, we train a 3D model utilizing a pretrained rectified flow model. The diffusion model’s trajectory is curved(Lipman et al., 2022), and the score (Song et al., 2020b; Poole et al., 2022) direction varies with different $t$. We denote $\epsilon_{\phi}(x_{t},t)$ as the score function, where $\phi$ represents the parameters of the denoise network, and $x_{t}$ follows the diffusion forward process. In contrast, the trajectory of the rectified flow model (Liu et al., 2022; 2023) is straight, and the vector field direction, $v_{\phi}(x_{t},t)$, remains the same for different $t$. In our VFDS framework – where $t\sim U[0,1]$ is used in every optimization step – the VFDS optimization direction is more consistent. However, the over-smoothing issue of SDS still exists. Therefore, we further analyze the grounding reasons for the over-smoothing issue from the perspective of ODE trajectories.
@@
When camera poses have only mild differences, the rendered images appear nearly identical. Different ODE trajectories cause inconsistent update directions, which means that directions of $\left(v_{\phi}(x_{t},t)-\left(\epsilon-x\right)\right)$ are inconsistent. Figure 4 illustrates a toy example. $\epsilon^{1},\epsilon^{2},\epsilon^{3}$ are noise randomly sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$ independently. According to Equation 1, we can get $x_{t}^{1},x_{t}^{2},x_{t}^{3}$, which looks like a blurred image (a hamburger with noise). Following the above analysis, the same $x$ (rightmost hamburger in Figure 4) together with different $\epsilon^{1},\epsilon^{2},\epsilon^{3}$ form different trajectories. During the VFDS training process, our 3D model takes $x_{t}^{1},x_{t}^{2},x_{t}^{3}$ as input and outputs the estimation of trajectory gradients (as shown in Figure 4). Note that the fitting targets are $\epsilon^{1}-x,\epsilon^{2}-x,\epsilon^{3}-x$ respectively. The model finally learns the multiple trajectories gradient, causing the over-smoothing issue.
@@ 3.2 Unique Couple Matching Loss
$x_{\delta_{T_{n-1}}}$ $=x_{\delta_{T_{n-1}}}+v_{\phi}(x_{\delta_{T_{n-2}}},\delta_{T_{n-2}})\delta_{T_{n-1}},n\geq 2$
where, $\sum_{i=0}^{n}\delta_{T_{i}}=1$, $\delta_{T_{0}}=0$. We define the process from $x$ to $\epsilon$ as push-backward process, denoted ##\##. And #$\#_{\phi}[x]$ denotes iteratively calculate $v_{\phi}(x_{\delta_{T_{i}}},t)$ to backtrack to the $\epsilon$ from $x$ in  Equation 6.
Due to the reversible property of rectified flow, we can search for a noise $\epsilon$ from $x$ in the VFDS framework (see Figure 3(b)). Additionally, because of the aforementioned coupling property, the search noise is unique.
By replacing randomly sampled noise $\epsilon$ to #$\#_{\phi}[x]$ in Equation 5, our UCM loss is defined as follows:
@@ 3.3 FlowDreamer for NeRF and 3D GS
Based on the UCM loss, we propose a novel framework, named FlowDreamer. Our framework yields high-fidelity results with richer textual details and faster convergence. FlowDreamer can be applied to two types of 3D models: 3D Gaussian splatting (Kerbl et al., 2023) and NeRF (Mildenhall et al., 2021).
Application to 3D Gaussian splatting. For 3D Gaussian splatting, we use points generated by the text-to-3D generator (Nichol et al., 2022) as parameter initialization. Then, we directly train the 3D Gaussian splatting model using our UCM loss. Our FlowDreamer yields high-fidelity results with richer textual details (see Figure 1). We also notice our method converges faster than VFDS (see Figure 9).
Application to NeRF and the initialization issue. For NeRF, however, a direct application does not perform well. We call this the initialization issue of NeRF, as we will explain the reasons below. When searching the noise $\epsilon$ for a given image $x$, i.e., push-backward process, we use the rectified flow model $v_{\phi}$. It defines a mapping from data distribution $\pi_{0}$ to noise distribution $\pi_{1}$, where $\pi_{0}$ is the distribution of its pre-training datasets. The effectiveness of the push-backward process depends on that the input distribution for $v_{\phi}$ should be aligned with or at least approximate to $\pi_{0}$. Otherwise, the input lies in an undefined area for $v_{\phi}$ hence the output is unreasonable.
When training NeRF from scratch, it can hardly generate reasonable images based on its randomly initialized parameters. Therefore, the input distribution (denoted as $\pi^{nf}$) is far from $\pi_{0}$, causing the rectified flow model $v_{\phi}$ difficult to estimate the gradient of the ODE trajectory. To solve this issue, we temporarily use the naive VFDS training to warm up as a remedy. We view this initialization issue as an open question and advocate further investigations.
(a) NeRF
(b) 3D Gaussian splatting
As for 3D GS models, the issue does not exist. Taking Figure 5 as an example, when the prompt “A English cottage with stone walls” is provided, the NeRF Initialization is simply a gray image, while the output of the 3D GS model (initialized by Point-E (Nichol et al., 2022)) resembles a cottage. This indicates the initial distribution of the 3D Gaussian splatting model is more approximate to $\pi_{0}$. The result of push-backward process is also more effective, as they are closer to the gradient from images output by the trained model.
@@ 4 Experiments
Figure 6: Qualitative comparison under 3D GS generation setting. Our FlowDreamer generates objects with finer details.
4.1 Implementation Details
We adopt Stable Diffusion 3 (SDv3) (Esser et al., 2024) as our Rectified flow model. To facilitate a better comparison, we will categorize the methods into two types: one where the 3D model is NeRF, and the other where the 3D model is 3D GS. For NeRF results comparisons, unlike ProlificDreamer(Wang et al., 2024a) and Consistent3D(Wu et al., 2024), which use multi-stage rendering with normal, geometric, and texture rendering, our method simply uses Instant-NGP (Müller et al., 2022) for rendering. It is optimized with a resolution of 256 for the first 5000 steps with a batch size of 1, and then 512 for the subsequent 3000 steps, also with a batch size of 1. For the warm-up strategy, we use SDS to optimize over 1200 steps. For 3D Gaussian splatting comparisons, we utilize the pretrained PointE (Nichol et al., 2022) to initialize the locations of 3D Gaussians, while other properties of 3D Gaussians adopt random initialization. The batch size and number of iterations are set to 4 and 3000, respectively.
4.2 3D Generation Comparison
Figure 7: Qualitative comparison under NeRF generation setting. Our FlowDreamer generates higher fidelity and sharper results.
NeRF Generation. We compare our FlowDreamer with DreamFusion (Poole et al., 2022), ProlificDreamer (Wang et al., 2024a), Consistent3D (Wu et al., 2024) in NeRF. Other SoTA baselines(Poole et al., 2022; Wang et al., 2024a; Wu et al., 2024) reimplemented by Three-studio(Guo et al., 2023) codebase and employ Stable Diffusion 2.1 for the prior.
As shown in Figure 7, our FlowDreamer achieves results with high fidelity and accurate text alignment. For example, the ‘saguaro cactus’ and ‘clay pot’ exhibit more detail and greater visual quality for the first prompt. Only FlowDreamer does not render the ‘octopus’ and ‘harp’ as a single object for the second prompt.
3D Gaussian splatting Generation. We compare our FlowDreamer with Dreamgaussian(Tang et al., 2023), Gaussiandreamer(Yi et al., 2023) and Luciddreamer(Liang et al., 2023). These 3D GS SoTA baselines are based on their official code by employing the Stable Diffusion 2.1 as the prior.
As shown in Figure 6, our method generates objects that match well with the input text prompts and exhibit realistic textures. For example, our generated ‘pumpkin’ is of high fidelity, and only our method generates the ‘spiders’, which match the text prompt ’plastic’ for the first prompt. The ‘origami pig’ has very rich details, such as its eyes and creases, which are relatively realistic for the second prompt. Our FlowDreamer shows an overall improvement in terms of visual quality and textural details.

### v3

@@ Abstract
Recent advances in text-to-3D generation have made significant progress. In particular, with the pretrained diffusion models, existing methods predominantly use Score Distillation Sampling (SDS) to train 3D models such as Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3D GS). However, a hurdle is that they often encounter difficulties with over-smoothing textures and over-saturating colors. The rectified flow model – which utilizes a simple ordinary differential equation (ODE) to represent a straight trajectory – shows promise as an alternative prior to text-to-3D generation. It learns a time-independent vector field, thereby reducing the ambiguity in 3D model update gradients that are calculated using time-dependent scores in the SDS framework. In light of this, we first develop a mathematical analysis to seamlessly integrate SDS with rectified flow model, paving the way for our initial framework known as Vector Field Distillation Sampling (VFDS). However, empirical findings indicate that VFDS still results in over-smoothing outcomes. Therefore, we analyze the grounding reasons for such a failure from the perspective of ODE trajectories. On top, we propose a novel framework, named FlowDreamer, which yields high-fidelity results with richer textual details and faster convergence. The key insight is to leverage the coupling and reversible properties of the rectified flow model to search for the corresponding noise, rather than using randomly sampled noise as in VFDS. Accordingly, we introduce a novel Unique Couple Matching (UCM) loss, which guides the 3D model to optimize along the same trajectory. Our FlowDreamer is superior in its flexibility to be applied to both NeRF and 3D GS. Extensive experiments demonstrate the high-fidelity outcomes and accelerated convergence of FlowDreamer. Moreover, we highlight the intriguing open questions, such as initialization challenges in NeRF and sampling techniques, to benefit the research community. For more details, please visit our project page at https://vlislab22.github.io/FlowDreamer.
@@ 1 Introduction
Recently, with the advances of text-to-2D image synthesis techniques based on the diffusion models, text-to-3D generation also undergoes a surge of research interest. A seminal work, DreamFusion (Poole et al., 2022) sets a cornerstone by proposing Score Distillation Sampling (SDS) to address this issue by leveraging pretrained text-to-image diffusion model, to train Neural Radiance Fields (NeRF) (Mildenhall et al., 2021). It has been rapidly evolved to 3D Gaussian Splatting (3D GS) (Kerbl et al., 2023; Tang et al., 2023; Yi et al., 2023) for faster training and rendering.
Despite the success, existing works  (Lin et al., 2023; Poole et al., 2022; Zhu et al., 2023) unveil that SDS suffers from issues such as over-smoothing textures and over-saturating colors. For this reason, some attempts, e.g.,Wang et al. (2024a); Liang et al. (2023); Wu et al. (2024) improve SDS from different perspectives. For example, ProlificDreamer (Wang et al., 2024a) introduces variational score distillation (VSD), which models 3D parameters as random variables to distill 3D assets. However, it requires significantly more time to optimize. Consistent3D (Wu et al., 2024) designs a consistency distillation sampling method to train the 3D model. Nevertheless, the quality improvements are still limited. LucidDreamer (Liang et al., 2023) proposes interval score matching (ISM) loss in the diffusing trajectory, but the loss is formulated based on strong assumptions and drops many terms with the same scale.
Recently, flow matching approaches (Liu et al., 2022; Lipman et al., 2022) pave new ways for fast and high-quality generation. Among them, rectified flow model (Liu et al., 2022; Esser et al., 2024; Liu et al., 2023) uses a simple ordinary differential equation (ODE) to represent a straight trajectory. It trains a time-independent vector field, indicating that the direction of the vector field remains constant. Whereas diffusion trajectories result in curved trajectories (Lipman et al., 2022). Score (Song et al., 2020b) is time-dependent, meaning that SDS optimizing over uniformly sampled values of $t$ can produce different gradient directions. Owning to its merits, we interestingly find that it could serve as an alternative prior for text-to-3D generation.
@@
In light of this, we first develop a mathematical analysis to seamlessly integrate SDS with rectified flow model. This enables us to build up an initial framework, named as Vector Field Distillation Sampling (VFDS). However, empirical results demonstrate that VFDS still leads to over-smoothing textures (See Figure 2). To this end, we further analyze the grounding reasons for such a failure from the perspective of ODE trajectories. This way, we find that, as VFDS randomly samples noise, it leads to multiple ODE trajectories in nearly the same images, i.e., from camera poses with mild differences (See Figure 5). This causes inconsistent update directions, leading to over-smoothing textural issues.
@@ 2 Related Works
Figure 3: Illustration of our FlowDreamer. Images of random views from different camera poses are sampled and then input to the VAE encoder to obtain the latents. We replace the randomly sampled noise $\epsilon$ in VFDS with #$\#_{\phi}[x]$ via the push-backward process. Next, we sample $t$ from $U[0,1]$ and interpolate to obtain $x_{t}$. Finally, the UCM loss with the conditional prompt is applied to update the 3D model.
@@ 3 FlowDreamer
Overview. An overview of our proposed FlowDreamer is depicted in Figure 3. The key insight is to leverage the coupling and reversible properties of the rectified flow model to search for the corresponding noise, rather than using randomly sampled noise as in our initial framework VFDS (see Sec. 3.1). Accordingly, we introduce a novel Unique Couple Matching (UCM) loss in Sec. 3.2, which guides the 3D model to optimize along the same trajectory. Finally, our FlowDreamer can be applied to two types of 3D models: 3D Gaussian splatting (Kerbl et al., 2023) and NeRF (Mildenhall et al., 2021) settings. Now let’s describe the details (see Sec. 3.3).
@@ 3.1 VFDS: SDS in the Lens of Rectified flow
Adapting SDS to the rectified flow framework. We first briefly introduce the rectified flow (Liu et al., 2022). Let $\pi_{1}$ and $\pi_{0}$ denote Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ and data distribution, respectively. $\epsilon$ and $x_{0}$ are respectively sampled from $\pi_{1}$ and $\pi_{0}$. Rectified flow defines the forward process as(to simplify the representation, below, $x_{0}$ and $x_{t}$ indicate the latent space.):
@@
where $w(t)$ is a time-dependent weighting function, $U[0,1]$ denotes the uniform distribution within $[0,1]$. Because the rectified flow model is an ODE model, it has reversible and coupling properties. The $\epsilon$ from the Gaussian noise distribution is uniquely coupled with the $x$ from the data distribution. Moreover, the rectified flow is reversible, as shown in Figure 4(a). Specifically,
@@
where $\theta$ is the 3D model parameters, $x=g(\mathbf{\theta,c})$ denotes a rendered image from a camera pose $\mathbf{c}$, $\epsilon$ denotes randomly sampled Gaussian noise, $t\sim U[0,1]$. Following the convention of the SDS, we omit the transformer Jacobian term for effective training. Therefore, $\left(\frac{\partial v_{\phi}(x_{t},t)}{\partial x_{t}}\frac{\partial x_{t}}{\partial x}+1\right)$ becomes a constant and can be absorbed by $w(t)$, so we have
@@
Figure 4: (a): An illustration of the reversible and coupling properties of the rectified flow model. The reversible property indicates that $\epsilon$ can map to $x_{0}$, and $x_{0}$ can map to $\epsilon$ by reversing the direction of $v_{\phi}$. The coupling property indicates that $\epsilon$ and $x_{0}$ can only form a unique coupling. For example, $\epsilon_{2}$ and $x_{0}^{2}$ form a coupling $(\epsilon_{2},x_{0}^{2})$; therefore, $\epsilon_{2}$ and $x_{0}^{1}$ can’t form a coupling $(\epsilon_{2},x_{0}^{1})$ again. (b): An illustration of the trajectories of diffusion and rectified flow. The gradient direction of the diffusion trajectory varies with different $t$, while the rectified flow roughly remains the same for different $t$ under ideal circumstances.
Based on Equation 5, we train a 3D model utilizing a pretrained rectified flow model. The diffusion model’s trajectory is curved (Lipman et al., 2022), and the score (Song et al., 2020b; Poole et al., 2022) direction varies with different $t$. (see Figure 4(b)). We denote $\epsilon_{\phi}(x_{t},t)$ as the score function, where $\phi$ represents the parameters of the denoise network, and $x_{t}$ follows the diffusion forward process. In contrast, the trajectory of the rectified flow model (Liu et al., 2022; 2023) is straight, and the vector field direction, $v_{\phi}(x_{t},t)$, roughly remains the same for different $t$ under ideal circumstances. In our VFDS framework – where $t\sim U[0,1]$ is used in every optimization step – the VFDS optimization direction is more consistent, therefore, its convergence speed is faster (please refer to the supplementary material). However, the over-smoothing issue of SDS still exists. Therefore, we further analyze the grounding reasons for the over-smoothing issue from the perspective of ODE trajectories.
@@
When camera poses have only mild differences, the rendered images appear nearly identical. Different ODE trajectories cause inconsistent update directions, which means that directions of $\left(v_{\phi}(x_{t},t)-\left(\epsilon-x\right)\right)$ are inconsistent. Figure 5 illustrates a toy example. $\epsilon_{1},\epsilon_{2},\epsilon_{3}$ are noise randomly sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$ independently. According to Equation 1, we can get $x_{t}^{1},x_{t}^{2},x_{t}^{3}$, which looks like a blurred image (a hamburger with noise). Following the above analysis, the same $x$ (rightmost hamburger in Figure 5) together with different $\epsilon_{1},\epsilon_{2},\epsilon_{3}$ form different trajectories. During the VFDS training process, the rectified flow model takes $x_{t}^{1},x_{t}^{2},x_{t}^{3}$ as input and outputs the estimation of trajectory gradients, as shown in Figure 5. Note that the fitting targets are $\epsilon_{1}-x,\epsilon_{2}-x,\epsilon_{3}-x$ respectively. The 3D model finally learns the multiple trajectories gradient, causing the over-smoothing issue.
@@ 3.2 Unique Couple Matching Loss
We define the process from $x$ to $\epsilon$ as push-backward process, denoted ##\##, which can be written as:
@@
$x_{\delta_{T_{n-1}}}$ $=x_{\delta_{T_{n-2}}}+v_{\phi}(x_{\delta_{T_{n-2}}},\delta_{T_{n-2}})\Delta_{T_{n-1}},n\geq 2$
where, $\sum_{i=1}^{n}\Delta_{T_{i}}=1$, $\delta_{T_{0}}=0$. And #$\#_{\phi}[x]$ denotes iteratively calculate $v_{\phi}(x_{\delta_{T_{i}}},t)$ to backtrack to the $\epsilon$ from $x$ in  Equation 6. Due to the reversible property of rectified flow, we can search for a noise $\epsilon$ from $x$ in the VFDS framework. Additionally, because of the aforementioned coupling property, the search noise is unique. By replacing randomly sampled noise $\epsilon$ to #$\#_{\phi}[x]$ in Equation 5, our UCM loss is defined as follows:
@@ 3.3 FlowDreamer for NeRF and 3D GS
Based on the UCM loss, we propose a novel FlowDreamer framework. It yields high-fidelity results with richer textual details and faster convergence. FlowDreamer can be applied to two types of 3D models: 3D Gaussian splatting (Kerbl et al., 2023) and NeRF (Mildenhall et al., 2021).
Application to 3D Gaussian Splatting. We use points generated by the text-to-3D generator (Nichol et al., 2022) as parameter initialization. Then, we directly train the 3D GS model using our UCM loss. Our FlowDreamer yields high-fidelity results with richer textual details (see Figure 1). It also turns out that our method converges faster than VFDS (please refer to the supplementary material).
Application to NeRF and the initialization issue. For NeRF, however, a direct application does not perform well. We call this the initialization issue of NeRF, as we will explain the reasons below. When searching the noise $\epsilon$ for a given image $x$, i.e., push-backward process, we use the rectified flow model $v_{\phi}$. It defines a mapping from data distribution $\pi_{0}$ to noise distribution $\pi_{1}$, where $\pi_{0}$ is the distribution of its pre-training datasets. The effectiveness of the push-backward process depends on that the input distribution for $v_{\phi}$ should be aligned with or at least approximate to $\pi_{0}$ or $\pi_{1}$. Otherwise, the input lies in an undefined area for $v_{\phi}$ hence the output is unreasonable.
When training NeRF from scratch, it can hardly generate reasonable images based on its randomly initialized parameters. Therefore, the input distribution (denoted as $\pi^{nf}$) is far from $\pi_{0}$ and $\pi_{1}$, causing the rectified flow model $v_{\phi}$ difficult to estimate the gradient of the ODE trajectory. To solve this issue, we temporarily use the naive VFDS training to warm up as a remedy. We view this initialization issue as an open question and advocate further investigations.
As for 3D GS models, the issue does not exist. The example can be found in the supplementary material, when the prompt “A English cottage with stone walls” is provided, the NeRF Initialization is simply a gray image, while the output of the 3D GS model (initialized by Point-E (Nichol et al., 2022)) resembles a cottage. This indicates the initial distribution of the 3D Gaussian splatting model is more approximate to $\pi_{0}$. The result of push-backward process is also more effective, as they are closer to the gradient from images output by the trained model.
@@ 4 Experiments
Figure 6: Qualitative comparison under 3D GS and NeRF generation setting. Our FlowDreamer generates objects with finer details.
4.1 3D Generation Settings
3D Gaussian Splatting Generation. We compare our FlowDreamer with DreamGaussian (Tang et al., 2023), GaussianDreamer (Yi et al., 2023) and LucidDreamer (Liang et al., 2023). These 3D GS SoTA baselines are based on their official code by employing the Stable Diffusion 2.1 as the prior. As shown in Figure 6(a), our method generates objects that match well with the input text prompts and exhibit realistic textures. For example, our generated ‘pumpkin’ is of high fidelity, and only our method generates the ‘spiders’, which matches the text prompt ’plastic’ for the first prompt. The ‘origami pig’ has rich details, such as its eyes and creases, which are relatively realistic for the second prompt. Although DreamGaussian (Tang et al., 2023) and GaussianDreamer (Yi et al., 2023) require comparatively less time, their results are generally subpar. Our FlowDreamer shows an overall improvement in terms of visual quality and textural details.
NeRF Generation. We compare our FlowDreamer with DreamFusion (Poole et al., 2022), ProlificDreamer (Wang et al., 2024a), Consistent3D (Wu et al., 2024) in NeRF. Other SoTA baselines (Poole et al., 2022; Wang et al., 2024a; Wu et al., 2024) reimplemented by Three-studio (Guo et al., 2023) codebase and employ Stable Diffusion 2.1 for the prior. As shown in Figure 6(b), our FlowDreamer achieves results with high fidelity and accurate text alignment. For example, the ‘saguaro cactus’ and ‘clay pot’ exhibit more detail and greater visual quality for the first prompt. Only FlowDreamer does not render the ‘octopus’ and ‘harp’ as a single object for the second prompt. Our FlowDreamer takes only more time than DreamFusion Poole et al. (2022), but the quality of DreamFusion’s results is limited. (For more results, please refer to the supplementary material.)